OpenStudy (anonymous):

8-10x-3x^2/x^2+9x+20 PLEASE SOLVE AND EXPLAIN

OpenStudy (jdoe0001):

have covered factoring of quadratics yet?

OpenStudy (anonymous):

yep

OpenStudy (jdoe0001):

.... ok.. one sec

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf \cfrac{8-10x-3x^2}{x^2+9x+20}\implies \cfrac{8-10x-3x^2}{(x+4)(x+5)}\) see how the denominator was factored?

OpenStudy (anonymous):

yes i do understand that. but how do i factor the numerator

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2&+9x&+20\\ &\uparrow &\uparrow \\ &4+5&4\cdot 5 \end{array}\) the sum of the components, add up to 9 their product gives 20

OpenStudy (anonymous):

but like the numerator is backward...how do i get the x^2 on the left side

OpenStudy (jdoe0001):

teh numerator is mainly trial and error you'd want to factor out the "8", the constant and factor out the "3", the leading term coefficient then fiddle and check which pair gives the middle and the last term so \(\large { \cfrac{8-10x-3x^2}{x^2+9x+20}\implies \cfrac{8-10x-3x^2}{(x+4)(x+5)} \\ \quad \\ \cfrac{-(3x^2+10x-8)}{(x+4)(x+5)}\quad \\ \quad \\ \begin{cases} 8=2\cdot 2\cdot \cdot 2\cdot 1\\ 3=3\cdot 1 \\\hline\\ (3\cdot +4)+(1\cdot -2)\implies +12-2\implies &10\\ -2\cdot +4\implies &-8\\ 3x\cdot x\implies &3x^2 \end{cases}\qquad thus \\ \quad \\ \cfrac{-(3x^2+10x-8)}{(x+4)(x+5)}\implies \cfrac{-[(3x-2)\cancel{(x+4)}]}{\cancel{(x+4)}(x+5)} }\)

OpenStudy (anonymous):

ok thank you

OpenStudy (jdoe0001):

yw